OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 2x - 3y - z if x = -2, y = 3, and z = -2

OpenStudy (anonymous):

2x - 3y - z (x = -2, y = 3, and z = -2) now we sub 2(-2) - 3(3) - (-2) can you do it now ?

OpenStudy (anonymous):

give it a try...what is 2 * -2 ?

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

- 4 - 9 + 2 = -13 + 2 = -11 sheelo is correct :)